OpenStudy (anonymous):

find the slope-intercept form of the equation of the line passing through the point (-3,2) and parallel to the line with the equation 2x+y=7

OpenStudy (anonymous):

You know what the slope intercept form is right?\[y = mx + b\]

OpenStudy (anonymous):

Isolate the "y" variable on the left side by moving your "x" variable onto the right side

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Did you get it? I can check for you :)

OpenStudy (anonymous):

Also, you're welcome!

OpenStudy (anonymous):

i'm still lost

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

yea, I get that, but then do I have to plug in (-3, 2) into the y= mx+b to find the other slope?

OpenStudy (anonymous):

Not the slope. What is the equation you have right now?

OpenStudy (anonymous):

we have to find the equation of the line parallel to 2x+y=7

OpenStudy (anonymous):

and it must go through (-3,2)

OpenStudy (anonymous):

I know that. But parallel lines have the same slope...

OpenStudy (anonymous):

right, so 2/1is the slope

OpenStudy (anonymous):

so you plug that in for m

OpenStudy (anonymous):

So here is the initial line\[ 2x+y=7 \]\[y=7 -2x\] From the equation I gave you, what is the slope? Not +2/1

OpenStudy (anonymous):

-2/1

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

CORRECT!

OpenStudy (anonymous):

so would the slope intercept form be\[y=-2x-4\]

OpenStudy (anonymous):

wait no!

OpenStudy (anonymous):

Since you're given a point that the line you're looking for will pass through, you can solve for b. Insert your points \[y = mx + b \]\[(+2) = -2(-3) + b\]\[2= 6 + b\]\[? = b \]

OpenStudy (anonymous):

\[-4\]

OpenStudy (anonymous):

Correct! Now plug in the results into your equation \[y = mx+b\]\[y = -2x -4\]So why did you doubt yourself?

OpenStudy (anonymous):

thanks so much!

OpenStudy (anonymous):

You're Welcome! :) Anytime